OpenStudy (anonymous):

Write and Solve Partial Fraction Decomposition of a Rational Expression (LOOK IN FOR EQUATION)

OpenStudy (anonymous):

\[\frac{ 2x }{ x ^{3}-1 }\]

OpenStudy (unklerhaukus):

how did you get your first step?

OpenStudy (unklerhaukus):

\[x^3-1=(x-1)(x^2+x+1)\]

OpenStudy (unklerhaukus):

so \[\frac{2x}{x^3-1}=\frac A{x-1}+\frac {Bx+C}{x^2+x+1}\]

OpenStudy (unklerhaukus):

the factor (x^2+x+1) is quadratic , so the denominator must be linear, not just a scalar

OpenStudy (unklerhaukus):

pardon?

OpenStudy (unklerhaukus):

in where?

OpenStudy (unklerhaukus):

im not sure what you are saying

OpenStudy (unklerhaukus):

from the start.

OpenStudy (unklerhaukus):

\[\frac{2x}{x^3-1}=\frac{2x}{(x-1)(x^2+x+1)}=\frac A{x-1}+\frac {Bx+C}{x^2+x+1}\] \[2x=A(x^2+x+1)+(Bx+C)(x-1)\] to find the constants try x=1 , x=0

OpenStudy (anonymous):

Is This Correct? b=-2/3 c=2 a=2/3

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]great work